• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Moffit

    Log entry time 12:47:04 on February 15, 2009

    Entry number 261297

    keyword=Trouble with Compton DAQ

    Having trouble getting the compton daq running.

    First problem was that the cavity power did not rise and fall while the DAQ was running. Tried a few CODA and ROC reboots.

    Now.. The photon rate has disappeared (EPICS reading 0).

    Paged Alexandre.